# Type 1 Diabetes Management using GLIMMER: Glucose Level Indicator Model with Modified Error Rate

### v1

@@ VII Conclusion
In this paper, we introduce GLIMMER, a machine learning algorithm with a custom loss function designed for accurate prediction of blood glucose levels and to create more reliable opportunities for behavioral and medical treatments in type 1 diabetes management. Our contribution emphasizes predictions in dysglycemic regions, where patients face dangerous conditions and require precise forecasts to prevent adverse events. Utilizing carefully selected input features and a custom loss function fine-tuned through a genetic algorithm, GLIMMER has demonstrated improved performance over state-of-the-art models, reducing RMSE by 23% and MAE by 31% on the OhioT1DM dataset. Additionally, we collected a new dataset containing CGM records and insulin delivery events from 25 patients with T1D, allowing us to validate GLIMMER’s generalizability on a larger, real-world dataset while also creating a valuable resource for further research. GLIMMER can be integrated into automated insulin delivery systems and smartphone applications, supporting patients and physicians in more accurately managing T1D and preventing dysglycemia.

### v2

@@ 8 Conclusion
In this paper, we introduce GLIMMER, a machine learning algorithm with a custom loss function designed to improve blood glucose predictions and create more reliable opportunities for behavioral and medical interventions in type 1 diabetes management. Our contribution focuses on accurate predictions in dysglycemic regions, where patients face the greatest risk and where existing models often underperform. By combining carefully selected input features with a loss function fine-tuned through a genetic algorithm, GLIMMER achieves substantial improvements over state-of-the-art models, reducing RMSE by 23% and MAE by 31% on the OhioT1DM dataset. We also collected and leveraged a new real-world dataset comprising CGM records and insulin delivery events from 25 patients with T1D, validating GLIMMER’s generalizability and creating a valuable resource for future research. These findings demonstrate that GLIMMER can be integrated into automated insulin delivery systems and smartphone applications, supporting patients and clinicians in managing T1D more effectively and preventing dangerous glycemic events.
9 Acknowledgment
This work was supported in part by the Mayo Clinic and Arizona State University Alliance for Health Care Collaborative Research Seed Grant Program under Award Number ARI-320598. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funding organization.